Circular No. 5664

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY ACTIVITY IN CYGNUS AND AQUILA S. Brandt, A. J. Castro-Tirado and N. Lund, Danish Space Research

Institute, report: “A new x-ray transient, EU 2053+316, has been discovered with the WATCH wide field monitor on the ESA EURECA satellite. First seen on Nov. 23, it has persisted since, but with varying intensity. The position (α = 20^{h}53^{m}, δ = +31^{o}.6, equinox 1950.0; estimated error radius 1^{o}) is inside the Cygnus Loop and near the bright northwestern rim as seen in the Einstein image of the Loop. Cyg X-2 flared on Nov. 14 and reached an intensity of 500 mCrab above 6 keV on Nov. 15, while Cyg X-3 flared to 300 mCrab on Nov. 17; both sources remain above the 100 mCrab level. The recurrent x-ray transient Aql X-1 (V1333 Aql) has been active since Nov. 20. GRS 1915+105 (IAUC 5590) has been monitored since the end of August and is still active and observed above 15 keV.”

PERIODIC COMET SWIFT-TUTTLE (1992t) G. Paubert, Institut de Radio Astronomie Millimétrique; P. Colom, D.

Bockelée-Morvan, J. Crovisier and L. Jorda, Observatoire de Paris-Meudon; and D. Despois, Observatoire de Bordeaux, report: “The millimeter spectrum of this comet was observed on Nov. 21 with the IRAM 30-m telescope. The following lines were detected (signal-to-noise ranging from 7 to 30), preliminary evaluations of the corresponding molecular production rates being given in units of 10^{27} s^{−1}: HCN J(1-0), 89 GHz, 0.25; CH_{3}OH, 12 lines around 145 and 165 GHz, 20; H_{2}S 1(1,0)-1(0,1), 168.8 GHz, 1.5; H_{2}CO 3(1,2)-2(1,1), 225.7 GHz, 2 (from a column density of 6 × 10^{12} cm^{−2} for a 13′′ HPBW and assuming a parent-molecule distribution). All line profiles are asymmetric with a cusp at negative velocities. The production rates of HCN and CH_{3}OH given on IAUC 5653 were underestimated due to ephemeris uncertainties.”

Jorda and J. Lecacheux, Paris-Meudon; and F. Colas, Observatoire de Paris, report: “Observations with the 1.05-m telescope and CCD camera at Pic du Midi during Nov. 20-26 showed a strong helicoidal jet extending from 1500 to 20 000 km. The jet rotates from one night to the next, and its motion is clearly seen between exposures taken over a 2-hr interval. Our observations are compatible with a nuclear rotation period of ∼ 2.9 days and an emission from a high-latitude active zone. From the observations in 1862 Z. Sekanina (1981, A.J. 86, 1741) deduced a period of 2.77 days.”

1992 November 27 Brian G. Marsden